# A Typology of Synthetic Datasets for Dialogue Processing in Clinical Contexts

### v1

@@ 3.1 A Typology for Human/Machine Microdata Interventions
  • Type 1. No intervention, i.e., there is no automatic or manual process of alteration, and data reflect naturally-occurring linguistic exchanges.
@@
  • Type 3. New microdata are produced de novo via a generative process of some kind, to substitute for “real” microdata. The generation may be conditioned on real data, and may also involve exogenously-provided knowledge in the form of prompts or templates. Iterative refinement involving human feedback or reinforcement learning (e.g., RLHF, human simulations) falls under this category when it results in the creation of fundamentally new data samples.
In our typology, these levels are not mutually exclusive. It is possible for a dataset to include both Type 2 and Type 3 interventions. Similarly, a dataset may involve multiple interventions of the same or a different type, focused on different aspects of its contents. Finally, we observe that synthesis interventions may be carried out by either humans or machines (or both) during the course of a dataset’s creation. In practice, many synthetic datasets involve a mix of both, and adequately capturing the ways in which a given dataset is “synthetic” requires that both aspects be described. As such, our typology involves categorizing a dataset along dimensions of both the human and machine involvement (Table 1).
Consider a dataset of “screen-written” patient-provider interactions, in which the dialogues’ contents and discursive structure are entirely imagined by humans in response to a prose case description (for example, the MTS-Dialog dataset (Ben Abacha et al., 2023)). As the representational goal of this dataset is to contain “realistic” clinical dialogue, to be used in place of otherwise unavailable genuine dialogues, we consider this to be a “Human Type 3, Machine Type 1” dataset.
@@
3.2 In Search of Synthetic Dialogues
To inform our typology’s design, we performed a literature review with the goal of surveying the landscape of synthetic biomedical dialogue datasets. We used keyword-based and indexed-term search strategies, supplemented by large language models (LLMs) to assist in filtering and classifying retrieved papers. See Appendix B for a detailed description of our approach and Appendix D for the verbatim LLM prompts. Table 2 summarizes the results of our structured search strategies across PubMed, the ACL Anthology, and dblp.
From a scientific standpoint, we note that clinical dialogue-oriented datasets are much more variable in terms of their linguistic domains, intended uses, structure, means of collection, methods of synthesis, etc. than the much larger but comparatively more homogeneous body of literature describing synthetic EHR notes. As such, they provide a highly informative corpus of examples to use in attempting to understand and categorize different approaches to synthesis. In practice, however, our typology (and this paper’s broader discussion of synthesis) would be generally applicable to datasets containing synthetic EHR notes or other forms of clinical language.
Given our focus on dialogues, however, we excluded from our search published datasets whose primary content was medical in nature but did not consist of dialogues, e.g., electronic health record (EHR) notes, structured EHR data, etc. This resulted in the exclusion of most existing work on data augmentation and synthetic medical data more generally, and on clinical language in particular. By far the most frequently-seen category of synthetic clinical language dataset is that of EHR notes.
We also excluded from our review datasets where the synthetic component was in the form of structured annotations (e.g. dialogue act classification, concept/entity labels, etc.), narrative summaries, or other secondary metadata, as opposed to the primary dialogue content itself. Additionally, we exclude papers that use genuine dialogues, but whose artifacts of interest and analysis are synthetically-produced secondary products (such as summaries or clinical notes ) generated automatically from those original genuine dialogues, .
Another category of dataset (and paper) that we encountered during our review described the creation of synthetic datasets that are not themselves presenting or simulating dialogues as such, but are intended to produce artifacts for use in training systems whose modality of use is conversational in nature. For example, Kang et al. (2024) describe a process for transforming curated biomedical data into pseudo-conversational question-answer pairs for use in instruction tuning a downstream LLM. While the intended purpose of such a system would be to support conversational interaction between a human and an LLM, the data itself are not conversational in nature nor does it attempt to characterize the key linguistic features of a dialogue-based interaction. As such, papers describing such datasets are of secondary interest and we have excluded them from our review. We note that, alongside synthetic EHR notes, synthetic QA pairs are also a very prevalent form of synthetic clinical dataset.
@@ 3.2 In Search of Synthetic Dialogues
For the purposes of our review, we defined our clinical domain of interest as comprising text capturing either spoken or written language generated by patients and/or providers, intended for use as part of clinical care, patient information-seeking, etc. As such, we do not take into account papers from the larger biomedical domain (scientific abstracts, descriptions of drug interactions, proteins, etc.).
@@
Another category that we ultimately excluded from our review are papers that generated synthetic dialogue data as a step towards some other primary analysis, but which did not release their resulting datasets as artifacts for use by others. One prototypical example of this category of work includes that of Yosef et al. (2024), who carried out a series of experiments designed to assess the ability of LLMs to simulate patient/therapist interactions. While they do produce synthetic dyadic interactions as part of their work, their primary interest is using them as inputs to another LLM-based analysis designed to determine the ability of LLMs to differentiate different styles and forms of therapy. A second example of this category of work would be that of Gray et al. (2024), who used LLMs to simulate patient-provider counseling discussions in a neonatology context and conducted a very thorough analysis in order to assess their content and utility from a clinical perspective.
@@
Classification 1: Clinical dialogue datasets classified according to different degrees of human-machine interventions.
Our final set of included papers may be found in Table 3. The classification shows an imbalance with fully-simulated (Type 3) data dominating the data creation methodologies. Additionally, with the advent of LLMs, an increasing trend of building machine-generated datasets is observable starting from the year 2023 onward.
4 Discussion
4.1 Synthetic Data and Varieties of Task
As discovered in our review, synthetic data sets in clinical domain are used for a variety of NLP related tasks such as: Information Extraction (Zhang et al., 2020), summarization of the clinical dialogues (Joshi et al., 2020; Chintagunta et al., 2021), generating notes from conversations Yim et al. (2020), dialogue with custom personality traits for training purposes Han et al. (2024), training data for ASR systems Fareez et al. (2022), and named entity recognition Frei and Kramer (2023).
Language Diversity in Datasets: As is commonly observed across NLP tasks, English was the most prominent language among the datasets we identified (e.g. Joshi et al. (2020); Chintagunta et al. (2021); Abercrombie and Rieser (2022); Wang et al. (2023a); Ben Abacha et al. (2023); Yim et al. (2023)). In addition, there is a growing amount of research for Chinese (Mandarin) data sets Zhang et al. (2020); Shi et al. (2023) as well as in Arabic ALMutairi et al. (2024), Korean (Han et al., 2024), and German (Röhrig et al., 2022; Frei and Kramer, 2023).
Type of Medical Domains: Most data sets we were able to retrieve are not targeted for a specific type of disease in the clinical domain, and instead attempted to more represent written clinical notes or patient-provider dialogues from generic clinical settings (e.g. an outpatient visit). However, we did encounter a handful of synthetic data sets focusing on specific clinical specialties, including cardiology (Zhang et al., 2020), gastroenterology (Liu et al., 2022), psychology (Liu et al., 2021; Zheng et al., 2023; Han et al., 2024), and obstetrics & gynecology (Wang et al., 2023b).
4.2 On Beyond Microdata
@@
There are also other factors that play important roles in the “realism” of a dataset. Recall the MTS-Dialog dataset of Ben Abacha et al. (2023), which consists of “dialogues” imagined and “screen-written” de novo by humans (in this case, trained clinicians) based on short clinical case vignettes. Now, imagine a hypothetical second dataset in which, instead of screen-writing a dialogue, the discourse samples were in fact the result of transcribed conversations taking place between two actors, posing as a doctor-patient dyad and improvisationally acting out a scenario based on the same case vignettes; the dataset described by Fareez et al. (2022) follows this method of creation, as does the aci subset of the Aci-bench dataset (Yim et al., 2023).
@@
In both of these examples, we would consider the resulting dataset to be synthetic, in the sense that neither one reflects the language produced by actual “real world” participants in the true interaction that the dataset is purporting to represent (i.e., a clinical encounter). Our typology in its present form would categorize both as Type 3 datasets, [^9] As the microdata are being “made up” (as opposed to having naturally arisen or being modified in situ). and we may safely say that the originating sociolinguistic context that gave rise to the linguistic information contained in both datasets is surely more artificial than in a hypothetical third dataset containing “genuine” conversations between patients and providers, captured “in the wild” (a Type 1 dataset, according to our typology).
@@
However, this thought experiment illustrates a limitation of our present typology. In this second scenario, the linguistic artifacts in the dataset (e.g., the utterances themselves, the paralinguistic and discursive content) were produced by actual humans in conversation, and could certainly be thought of as being “less synthetic” than the utterances in MTS-Dialog, which were made up from whole cloth. In both scenarios, the semantic content of the utterances may be quite similar, as might the degree to which they accurately reflect the vignettes that formed their seeds. What would vary would be the above-mentioned pragmatic and discursive aspects of the data, as a side effect of the difference in the method of generation. Our typology currently lacks a notion of (for lack of a better word) modality (imagined vs. transcribed spoken language). It also does not distinguish between degrees of synthesis along these dimensions.
@@
Another area that our typology does not capture is what we might think of the relationship between synthesis and a dataset’s sociolinguistic context. Consider a hypothetical dataset of transcribed English-language patient-provider conversations taking place during genuine outpatient visits in the United States (i.e., a Type 1, “real” dataset in our typology). Imagine, now, that the dataset were run through a machine translation (MT) algorithm, with the intent of producing a French-language dataset for use in evaluating algorithms designed to process Francophone clinical encounters. According to our typology, this is now a “Type 2” dataset: the dataset’s microdata are still clearly derived from “real” microdata, but have been perturbed in some systematic way.
@@
Depending on the quality of the machine translation system, the resulting dataset may be “equivalent” in terms of its lexical and syntactic representativeness, and may also be equivalently useful in terms of its semantic content (i.e., the set of clinical conditions and events being described). [^10] We note that the extent to which this is true will depend heavily on the specifics of its intended environment of use. If our hypothetical French-speaking clinical population is in France itself, the difference may be relatively small. If it is meant for use elsewhere in the Francophone world, there may be specific medical concepts and events that are missing or over-represented. In such a situation, we would argue that the translated dataset was of limited external validity. In other words, the impact of the synthesis may be minimal along those dimensions. However, the pragmatic and discursive content may be wildly inappropriate, as norms of clinical communication and patient expectations vary significantly between cultures and countries, and French providers and patients will not speak the same way to one another as their American counterparts would. [^11] We note that, even within societies, norms of clinical communication vary greatly across numerous social and demographic factors, as well as between clinical specialties. This is one of the reasons why we are particularly wary of synthetic data in the context of clinical dialogues. As such, even assuming a very high-quality translation, the dataset is likely no longer as valid, in terms of its linguistic representativeness (as it may have been before being made synthetic). This is likely the case to a degree that is much larger than might be expected by a linguistically naïve researcher equipped with an MT algorithm. [^12] Despite technical advances, translation remains, of course, notoriously challenging (Anonymous, 1968) and finicky (Carlson et al., 2007). Clinical contexts bring numerous specialized challenges for the use of MT (Mehandru et al., 2022).
@@
Additionally, consider that the extra-linguistic factors represented in our hypothetical dataset may be severely impacted by this sort of perturbation. In this particular example, the translated conversation may indeed be adequately representative in terms of its linguistic and clinical content, but its contextual representativeness may be impaired. A patient-provider discussion about the management of end-stage kidney disease will include discussion of lab values, diet, etc., all of which may translate into a French context with adequate fidelity. However, the parts of the conversation involving the logistics of insurance coverage for dialysis treatments, or the details of organizing a social worker to assist in transportation to and from dialysis sessions, would differ wildly between American and French contexts.
@@
A similar case exists in our data, where Aci-bench (a Human Synthetic Type 3 dataset) is used in a one-shot setting to prompt LLMs to generate equivalent dialogues in the Najdi Arabic dialect, ensuring cultural adaptation ALMutairi et al. (2024). According to our typology, the resulting data is machine-generated constrained by human-synthetic and falls under Human Type 3, Machine Type 3. As LLMs are increasingly deployed at scale, hybrid synthetic datasets of this type, will likely become more common. However, the validity and representativeness of such datasets must be carefully assessed.
5 Conclusion
@@
Academic literature on synthetic data focuses heavily on the mechanics of how synthetic data elements are generated, and our typology above focuses on the role that the synthesis plays in the overall “shape” of the dataset. Given that Type 3 synthetic human- and (especially) machine-generated clinical dialogue data are becoming increasingly prevalent, our work offers a broader perspective on syntheticity, helping researchers better assess and contextualize the validity and representativeness of these datasets.
The relationship between synthesis and a dataset’s larger contextual factors remains currently unexplored. As future work, we will expand our work to capture realism and synthesis along non-linguistic and contextual factors. [^13] The extent to which there exists a valid a dichotomy between “linguistic” and “non-linguistic” content in a dataset is surely up for debate.
6 Limitations

### v2

@@ 3.1.  A Typology for Human/Machine Microdata Interventions
  • Type 1. No intervention (i.e., there is no automatic or manual process of alteration, and data reflect naturally-occurring linguistic exchanges).
@@
  • Type 3. New microdata are produced de novo via a generative process of some kind, to substitute for “real” microdata. The generation may be conditioned on real data, and may also involve exogenously-provided knowledge in the form of prompts or templates. Iterative refinement involving human feedback or reinforcement learning (e.g., via RLHF [^8] Reinforcement Learning from Human Feedback (christianoEtAl2017RLHF). or human simulations) falls under this category when it results in the creation of fundamentally new data samples.
In our typology, these levels are not mutually exclusive. It is possible for a dataset to include both Type 2 and Type 3 interventions. Similarly, a dataset may involve multiple interventions of the same or a different type, focused on different aspects of its contents.
Consider a dataset of “screen-written” patient-provider interactions, in which the dialogues’ contents and discursive structure are entirely imagined by humans in response to a prose case description (for example, the MTS-Dialog dataset (ben-abacha-etal-2023-empirical)). As the representational goal of this dataset is to contain “realistic” clinical dialogue, to be used in place of otherwise unavailable genuine dialogues, we consider this to be a “Human Type 3, Machine Type 1” dataset.
@@
3.2.  In Search of Synthetic Dialogues
@@
Our typology’s design was informed by a literature review whose goal was to survey the landscape of synthetic biomedical dialogue datasets. We defined our clinical domain of interest as comprising text capturing either spoken or written language generated by patients and/or health providers, intended for use as part of clinical care, patient information-seeking, etc. Our goal was to identify published datasets of clinical dialogues, in spoken or written genre, and either synthetic or natural in nature.
Included publications could be specifically focused on the production and evaluation of their datasets, or could produce datasets as secondary research artifacts. In the latter case, we required that there be an intention on the part of the authors that their datasets be re-used by others (i.e., their data were not simply made re-usable for replicability). We used keyword-based and index-term search strategies, supplemented by large language models (LLMs) to assist in filtering and classifying retrieved papers.
From a total of 1,626 papers (679 from PubMed, 591 from ACL Anthology, and 356 from dblp), successive filtering steps (including LLM-assisted classification and manual review) reduced the pool to 25 included papers (4 from PubMed, 11 from ACL Anthology, and 10 from dblp). After removing overlapping entries, we arrived at 20 unique papers.
From a scientific standpoint, we note that clinical dialogue-oriented datasets are more variable in terms of their linguistic domains, intended uses, structure, means of collection, methods of synthesis, etc. than the larger but comparatively more homogeneous body of literature describing synthetic electronic health record (EHR) notes. As such, they provide an informative corpus of examples to use in attempting to understand and categorize different approaches to synthesis. In practice, however, our typology (and this paper’s broader discussion of synthesis) would be applicable to datasets containing synthetic notes or other forms of clinical language.
Given our focus on dialogues, we excluded the following data sets from our analyses in this paper:
1. Published datasets whose primary content was medical in nature but did not consist of dialogues, e.g., electronic health record (EHR) notes, structured EHR data, etc. This resulted in the exclusion of most existing work on data augmentation and synthetic medical data more generally, and on clinical language in particular. By far the most frequently-seen category of synthetic clinical language dataset is that of EHR notes.
2. Datasets where the synthetic component was in the form of structured annotations (e.g., dialogue act classification, concept/entity labels, etc.), narrative summaries, or other secondary metadata, as opposed to the primary dialogue itself.
3. Datasets that use genuine dialogues, but whose artifacts of interest and analysis are synthetically-produced secondary products (such as summaries or clinical notes ) generated automatically from those original genuine dialogues, .
We also excluded datasets that are not themselves presenting or simulating dialogues as such, but which are intended to produce artifacts for use in training systems whose modality of use is conversational in nature. For example, kang-etal-2024-self describe a process for transforming curated biomedical data into pseudo-conversational question-answer pairs for use in instruction tuning a downstream LLM. While the intended purpose of such a system would be to support conversational interaction between a human and an LLM, the data itself are not conversational in nature nor does it attempt to characterize the key linguistic features of a dialogue-based interaction.
As such, papers describing such datasets were of secondary interest, and we have excluded them from our review. We note that, alongside synthetic EHR notes, synthetic QA pairs are also a very prevalent form of synthetic clinical dataset.
@@
Another category that we ultimately excluded from our review are papers that generated synthetic dialogue data as a step towards some other primary analysis, but which did not release their resulting datasets as artifacts for use by others. One prototypical example of this category of work includes that of yosef2024assessing, who carried out a series of experiments designed to assess the ability of LLMs to simulate patient/therapist interactions. While they do produce synthetic dyadic interactions as part of their work, their primary interest is using them as inputs to another LLM-based analysis designed to determine the ability of LLMs to differentiate between styles and forms of therapy. A second example of this category of work would be that of grayIncreasingRealismVariety2024, who used LLMs to simulate patient-provider counseling discussions in a neonatology context and conducted a very thorough analysis in order to assess their content and utility from a clinical perspective.
@@
Figure 1: A selection of clinical dialogue datasets classified according to different degrees of human-machine interventions.
4.  Discussion
4.1.  Synthetic Data and Varieties of Task
As discovered in our review, synthetic data sets in clinical domain are used for a variety of NLP related tasks such as: Information Extraction (zhang-etal-2020-mie), summarization of the clinical dialogues (joshi-etal-2020-dr; chintagunta-etal-2021-medically), generating notes from conversations yim-etal-2020-alignment, dialogue with custom personality traits for training purposes han-etal-2024-psydial-personality, training data for ASR systems fareez2022dataset, and named entity recognition frei2023annotated.
Language Diversity in Datasets: As is commonly observed across NLP tasks, English was the most prominent language among the datasets we identified (e.g. joshi-etal-2020-dr; chintagunta-etal-2021-medically; abercrombie-rieser-2022-risk; wang-etal-2023-umass; ben-abacha-etal-2023-empirical; yim2023aci). In addition, there is a growing amount of research for Chinese (Mandarin) data sets zhang-etal-2020-mie; shi-etal-2023-midmed as well as in Arabic almutairi-etal-2024-synthetic, Korean (han-etal-2024-psydial-personality), and German (rohrig2022grascco; frei2023annotated).
Type of Medical Domains: Most data sets we were able to retrieve are not targeted for a specific type of disease in the clinical domain, and instead attempted to more represent written clinical notes or patient-provider dialogues from generic clinical settings (e.g. an outpatient visit). However, we did encounter a handful of synthetic data sets focusing on specific clinical specialties, including cardiology (zhang-etal-2020-mie), gastroenterology (liu2022meddg), psychology (liu-etal-2021-towards; zheng-etal-2023-augesc; han-etal-2024-psydial-personality), and obstetrics & gynecology (Wang:2023d).
4.2.  On Beyond Microdata
@@
There are also other factors that play important roles in the “realism” of a dataset. Recall the MTS-Dialog dataset of ben-abacha-etal-2023-empirical, which consists of “dialogues” that were imagined and “screen-written” de novo by humans (in this case, trained clinicians) based on short clinical case vignettes. Now, imagine a hypothetical second dataset (“dataset B”) in which, instead of being screen-writiten dialogue, the discourse samples were in fact the result of transcribed conversations taking place between two actors, posing as a doctor-patient dyad and improvisationally acting out a scenario based on the same case vignettes; the dataset described by fareez2022dataset follows this method of creation, as does the aci subset of the Aci-bench dataset (yim2023aci).
@@
In both of these examples, we would consider the resulting dataset to be synthetic, in the sense that neither one reflects the language produced by actual “real world” participants in the true interaction that the dataset is purporting to represent (i.e., a clinical encounter). Our typology in its present form would categorize both as Type 3 datasets along their human dimensinos, [^9] As the microdata are being “made up” (as opposed to having naturally arisen or being modified in situ). and in both cases we can safely say that the context that gave rise to the linguistic information contained in the dataset is surely more artificial than in a hypothetical third dataset (“dataset C”) containing “genuine” conversations between patients and providers captured “in the wild” (a Type 1 dataset, according to our typology).
@@
However, this thought experiment illustrates a limitation of our present typology. The linguistic artifacts in “dataset B” (e.g., the utterances themselves, along with their associated paralinguistic and discursive content) were produced by actual humans in conversation, and could certainly be thought of as being “less synthetic” than the utterances in MTS-Dialog, which were made up from whole cloth. In both scenarios, the semantic content of the utterances may be quite similar, as might the degree to which they accurately reflect the vignettes that formed their seeds. The variation would involve the above-mentioned pragmatic and discursive aspects of the data, as a side effect of the difference in the method of generation. Our typology currently lacks a notion of, for lack of a better word, modality: in this case, “imagined” vs. “transcribed” spoken language. It also does not distinguish between degrees of synthesis along these dimensions.
@@
Another area that our typology does not capture is the relationship between synthesis and a dataset’s sociolinguistic context nguyen-etal-2016-survey. Consider a hypothetical dataset of transcribed English-language patient-provider conversations taking place during genuine outpatient visits in the United States (i.e., a Type 1, “real” dataset in our typology). Imagine, now, that the dataset were run through a machine translation (MT) algorithm, with the intent of producing a French-language dataset for use in evaluating algorithms designed to process Francophone clinical encounters. According to our typology, this is now a “Type 2” dataset: the dataset’s microdata are still clearly derived from “real” microdata, but have been perturbed in some systematic way.
@@
Depending on the quality of the machine translation system, the resulting dataset may be “equivalent” in terms of its lexical and syntactic representativeness, and may also be equivalently useful in terms of its semantic content (i.e., the set of clinical conditions and events being described). [^10] We note that the extent to which this is true will depend heavily on the specifics of its intended environment of use. If our hypothetical French-speaking clinical population is in France itself, the difference may be relatively small. If it is meant for use elsewhere in the Francophone world, there may be specific medical concepts and events that are missing or over-represented. In such a situation, we would argue that the translated dataset was of limited external validity. In other words, the impact of the synthesis may be minimal along those dimensions. However, the pragmatic and discursive content may be wildly inappropriate, as norms of clinical communication and patient expectations vary significantly between cultures and countries, and French providers and patients will not speak the same way to one another as their American counterparts would. [^11] We note that, even within societies, norms of clinical communication vary greatly across numerous social and demographic factors, as well as between clinical specialties. This is one of the reasons why we are particularly wary of synthetic data in the context of clinical dialogues. As such, even assuming a very high-quality translation, the dataset is likely no longer as valid, in terms of its linguistic representativeness, as it may have been before being made synthetic. This is likely the case to a degree that is much larger than might be expected by a linguistically naïve researcher equipped with an MT algorithm. [^12] Despite technical advances, translation remains, of course, notoriously challenging (Anonymous:1968p9217) and finicky (Carlson:2007p3777). Clinical contexts bring numerous specialized challenges for the use of MT (mehandru_2022).
@@
Additionally, consider that the extra-linguistic factors represented in our hypothetical dataset may be severely impacted by this sort of perturbation. In this particular example, the translated conversation may indeed be adequately representative in terms of its linguistic and clinical content, but its contextual representativeness may be impaired. A patient-provider discussion about the management of end-stage kidney disease will include discussion of lab values, diet, etc., all of which may translate into a French context with adequate fidelity. However, the parts of the conversation involving the logistics of insurance coverage for dialysis, or of organizing a social worker to assist in transportation to and from dialysis sessions, would differ wildly between American and French contexts.
@@
This vignette is inspired by an actual example we encountered during our literature review, in which Aci-bench (a Human Synthetic Type 3 dataset) was used in a one-shot setting to prompt LLMs to generate equivalent dialogues in the Najdi Arabic dialect, with the goal of creating a dataset for use in cultural adaptation of an LLM almutairi-etal-2024-synthetic. According to our typology, the resulting data is machine-generated constrained by human-synthetic and falls under Human Type 3, Machine Type 3. As LLMs are increasingly deployed at scale, hybrid synthetic datasets of this type, will likely become more common. However, the validity and representativeness of such datasets must be carefully assessed.
5.  Conclusion
@@
Academic literature on synthetic data focuses heavily on the mechanics of how synthetic data elements are generated, and our typology focuses instead on the role that the synthesis plays in the overall “shape” of the dataset. Our typology provides authors with the ability to speak in a more clear and granular way about what aspects of their dataset are syntheetic, to what degree, and from what source (human vs. machine).
Our literature review revealed a clear pattern of increased prevalence of fully-synthetic (Type 3) datasets, particularly of machine-generated dialogues following the widespread availability of LLMs in 2023. Given this, our works’s broader perspective on syntheticity will help researchers better assess and contextualize the validity and representativeness of these datasets.
The relationship between synthesis and a dataset’s larger contextual factors remains currently unexplored. As future work, we hope to expand our work to capture realism and synthesis along non-linguistic and contextual factors. [^13] The extent to which there exists a valid a dichotomy between “linguistic” and “non-linguistic” content in a dataset is surely up for debate.
6.  Limitations